OpenStudy (geekfromthefutur):

Help me with this last question I really need help i give medal and i will fan

OpenStudy (ibbutibbu.):

thats a hard question man

OpenStudy (anonymous):

omg

OpenStudy (anonymous):

why did ibbutibbu put that ?

OpenStudy (anonymous):

can you put in the equation

OpenStudy (anonymous):

ehh

OpenStudy (geekfromthefutur):

@ganeshie8

OpenStudy (anonymous):

i cant does this

OpenStudy (geekfromthefutur):

@Compassionate @mathmale

OpenStudy (anonymous):

@driftracer305 please help them

OpenStudy (mathmale):

Any "solution" will be a point, a point that's on the straight line given by 5x-y=8. Go ahead and find one such point. Arbitrarily let x=1. Find y. Write these two coordinates in the form (x,y).

OpenStudy (geekfromthefutur):

im confused

OpenStudy (driftracer305):

there can be a number of solutions............ so u can just plugging with any numbers u like in part A...... start with 1 and keep going @geekfromthefutur

OpenStudy (mathmale):

You are given 5x-y=8. Replace that x by '1'. What do you get?

OpenStudy (mathmale):

My '1' is such an "any number u like.'

OpenStudy (geekfromthefutur):

51-7=8

OpenStudy (mathmale):

Excuse me, I should have asked you to "replace x with (1)." Please do that now. Then simplify 5(1).

OpenStudy (driftracer305):

5-8 = y = y = -3

OpenStudy (geekfromthefutur):

5-8=-3

OpenStudy (mathmale):

@driftracer305: If you look carefully, I'm trying to help our friend geek to solve this problem himself or herself. Kindly do NOT give away answers without the involvement of the person asking the question!

OpenStudy (mathmale):

'geek': Yes. if x=1, y=-3. Write that point in the form (x,y) now, please.

OpenStudy (geekfromthefutur):

OHHHHHH i think i get it

OpenStudy (driftracer305):

k @mathmale ........... sorry man.......

OpenStudy (mathmale):

You may then choose a couple of other x values and find the associated y-values. That takes care of Part B. Use material from our discussion to answer Part A. Are you OK with Part C?

OpenStudy (geekfromthefutur):

No i still dont get it the confusion is coming from the write the point in the form is wat is makin me stuck

OpenStudy (mathmale):

All points are written in the form (x,y). The x represents the 'x-coordinate" of the point. The y repres. the "y-coordinate" of the point. If, as in this example, x=1 and y=-3, then the point is (1,-3). choose 2 other x values. Find the corresponding y values. Than write the 2 points in the form (x,y), please.

OpenStudy (geekfromthefutur):

Oh

OpenStudy (geekfromthefutur):

So i can chose any number

OpenStudy (geekfromthefutur):

for the x value

OpenStudy (geekfromthefutur):

(2,-3) (3,-3)

OpenStudy (mathmale):

Toss a die and use whatever number comes out on top as your x-value. Really: ANY number is fine.

OpenStudy (mathmale):

If you choose x=2, then 5(x)-y=8 becomes 5(2)-y=8, or y=2. Check that. Do you agree or disagree? If you disagree, re-write the first point as (2, ? )

OpenStudy (geekfromthefutur):

i agree i suck at math somtimes

OpenStudy (mathmale):

Kindly don't say that...the least you can do for yourself is to say, "I can learn this just fine with a sufficient investment of time and effort." If x=2, then y=2. Thus, 5(2)-2=8 becomes true. Show how you got your 3rd point. Let x= 3. Where does that -3 come from? i'd like to see your work.

OpenStudy (geekfromthefutur):

i picked a random number

OpenStudy (mathmale):

You picked a random x-value, or you picked a random y-value? Pick a random x-value (NOT y-value) and use the given equation to calculate y.

OpenStudy (geekfromthefutur):

x value i kept the y the same as the others

OpenStudy (mathmale):

No. that equation governs the relationship between x and y. Given x, you CALCULATE y. or.... Given y, you CALCULATE x. Choose x=5. Substitute that into 5x-y=8. Calculate y, please.

OpenStudy (geekfromthefutur):

-3

OpenStudy (mathmale):

When I say "substitute that into 5x-y=8, I mean the following: Rewrite 5x-y=8 as 5(5)-y=8. What is 5(5)?

OpenStudy (geekfromthefutur):

5*5=25

OpenStudy (geekfromthefutur):

or 1

OpenStudy (mathmale):

Right. Then your 5x-y=8 becomes 25-y=8. Please now subtract 8 from both sides. What do you get?

OpenStudy (geekfromthefutur):

-25

OpenStudy (mathmale):

Simplify: 25-8-y=8-8

OpenStudy (geekfromthefutur):

0 or 9

OpenStudy (mathmale):

Have to write the whole equation. 25-8 is 17. 25-8-y is 17-y. Right side: 8-8 is zero. Write the left side (from above) , followed by the '=' sign, and then the right side (from above).

OpenStudy (geekfromthefutur):

So it 25-8 = 8-8

OpenStudy (mathmale):

17-y=0, actually. Make certain you understand where that came from. 25-8 is 17. 25-8-y is 17-y.

OpenStudy (mathmale):

You've dropped the 'y'. Can't do that.

OpenStudy (geekfromthefutur):

o ok

OpenStudy (mathmale):

25-8-y=8-8 becomes 17 - y = 0 Solve this for y, please.

OpenStudy (geekfromthefutur):

y=17

OpenStudy (mathmale):

Right. So, if x=5, y=17. Are you now OK with this, or have y ou questions about tghe process?

OpenStudy (geekfromthefutur):

No i good

OpenStudy (mathmale):

OK. So now you have 3 points that line on the line represented by 5x-y=8. OK? You should list them as you answer this question. I need to get off the 'Net pretty quickly. But do you have further questions about Parts A, B or C?

OpenStudy (geekfromthefutur):

So for part a it is 5-8=-3 part b is (1,-3) 2,-3 and 5,-3 part c is 17-y=0 and y=17

OpenStudy (geekfromthefutur):

Im just nailing it down to the answers

OpenStudy (mathmale):

Part A: You need to represent ALL possible solutions of 5x-y=8. That equation alone does that. If you want to be fancier, you could solve that equation and obtain y=5x-8. I am not going to check your results for Part B since we've discussed them in detail and you have our conversation to use to check your own results. What about Part C?

OpenStudy (geekfromthefutur):

3 points that line on the line represented by 5x-y=8.

OpenStudy (mathmale):

Yes...that's what you now have for Part B. Do you hve to do Part C? If so, what are your thoughts on that so far?

OpenStudy (geekfromthefutur):

part c is (17,5)

OpenStudy (mathmale):

No, sorry. You are to write an EQUATION similar to 5x-y=8 such that the two lines represented by these two equations never intersect.

OpenStudy (geekfromthefutur):

25-8-y=8-8

OpenStudy (mathmale):

"inconsistent" means / signifies that the two lines represented by your two different equations never intersect. Your 25-8-y=8-8 is the same equation you started with, and thus could not be the answer. Given 5x-y=8, you can obtain a new line that never intersects the given line simply by changeing the constant on the right. Do that now, please.

OpenStudy (mathmale):

Start with 5x-y=8. Throw out the 8 and choose some other number.

OpenStudy (geekfromthefutur):

17- y=0

OpenStudy (mathmale):

All you have to do is to modify the equation 5x-y=8 so that the number on the right is differnt from 8. Think of a number, any number, different from 8. What is it?

OpenStudy (geekfromthefutur):

5x-y=4

OpenStudy (mathmale):

Fine. Since 4 is different from 8, you have the equation of a line parallel to 5x-y=8. These two lines will never intersect. Thus, your system consissting of these two lines is "inconsistent." You're done.

OpenStudy (geekfromthefutur):

Thx for you time mathmale to help me on this

OpenStudy (mathmale):

My pleasure. Later!! Bye.